Two masses of 50.0 grams and 100 grams are placed on a 200-gram uniform meter stick. The 50 g mass is placed on the 10.0 cm mark while the 100 g is at 80.0 cm mark. Locate the center of gravity of the system of weights including the meterstick.
1
Expert's answer
2021-10-28T17:33:45-0400

We know that


"50\times(70-10)+200(70-x)=100(80-70)"

"50\times60+200\times(70-x)=100\times10"

"3000-1000=-200(70-x)"

"x=60cm"